Your company sponsors a 401(k) plan into which you deposit 14 percent of your $65,000 annual income. Your company matches 50 percent of the first 5 percent of your earnings. You expect the fund to yield 8 percent next year. Assume you are currently in the 31 percent tax bracket.

a, What is the total annual investment in the 401(k) plan at year-end? $Answer 1

b, What is your one-year return?
Answer #1

annual income = $65000

employee deposit = 14%

= 65000*14%

= 9100

tax rate given = 31%

tax savings = 9100*31%

= 2821

net employee cost = 9100 - 2821

= 6279

employer deposit = 65,000 * 50% * 5%

= 1625

Total employee and employer investment = 9100 + 1625

= 10725

given yield rate = 8%

a)

total investment at end of year 1 = 10725*(1+yield)

= 10725*1.08

= 11583

b)

one year return = (year end investment - employee net cost) / employee net cost

= (11583 - 6279) / 6279

= 0.8447 or 84.47%